.ÀŁþŊüŁõ ŒŁŮ cüŁĺďŚŁê ýŚŁûüŁþŚŁœŘĄŁŮ Ç c ŁþŚŁÞŁœ ď cĄŁÑŁĂŁõ ĆŁŞ üŁőŚŁþď ýùćŚŁĺ ďŚŁČŁÆŁŞ ñ cĄŁõÂŁê ŒŁþ ÂŁþě ďć

ÀŁŃŁŃŁŃŁŃŁŃŁ§_{x→π/10} sin x = \frac{1}{4} \sqrt{5} − 1

The above text is a very simple demonstration of FarsiTEX’s ability to typeset Persian mathematical formulas. ?ćÂŁŁŁŁŁŁŁŁŁŁŁŁî ŘÂŁŁŁŁŁŁŁŁŁŁŁŁŻŘ ÃŁŁŁŁŁŁŁŁŁŁŁŁČŁŁŁŁŁŁŁŁŁŁŁŁœ (Linux) ÅŁŁŁŁŁŁŁŁŁŁŁŁØŁŁŁŁŁŁŁŁŁŁŁŁĂ cŁŁŁŁŁŁŁŁŁŁŁŁČŁŁŁŁŁŁŁŁŁŁŁŁó ťc ŁŁŁŁŁŁŁŁŁŁŁŁ¡ŁŁŁŁŁŁŁŁŁŁŁŁŮ Řď ŒŁŮ cüŁŁŁŁŁŁŁŁŁŁŁŁĺďŚŁŁŁŁŁŁŁŁŁŁŁŁê öŘĄŁŁŁŁŁŁŁŁŁŁŁŁŮüŁŁŁŁŁŁŁŁŁŁŁŁõ ŚŁŁŁŁŁŁŁŁŁŁŁŁþŊ .ôĚŁŁŁŁŁŁŁŁŁŁŁŁĺ

?ćÂŁî ŘÂŁŻŘ ÃŁČŁœ (Linux) ÅŁØŁĂ cŁČŁó ťc Ł¡ŁŮ Řď ŒŁŮ cüŁĺďŚŁê öŘĄŁŮüŁõ ŚŁþŊ .ôĚŁĺ

And the above two lines (try to) demonstrate the Persian specific Line Justification method, implemented in FarsiTEX.